OpenStudy (anonymous):

Which polynomial is a difference of squares? x2 – 16x + 64 x2 + 4 9x2 – x – 36 25x2 – 36

OpenStudy (anonymous):

x^2-16x+64 = (x-8)(x-8) x^2+4 =x^2+2^2 i am not sure i got your q.

OpenStudy (anonymous):

it ok i got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

heres a medal